OpenStudy (prince1342):

Which of the following is a factor of 2x4 + 22x3 + 60x2?

OpenStudy (mathstudent55):

What common factor do the numbers 2, 22, and 60 have?

OpenStudy (mathstudent55):

Here are the prime factors of each of the three numbers above: \(2 = \color{red}{2}\) \(22 = \color{red}{2} \times 11\) \(60 = \color{red}{2} \times 2 \times 3 \times 5\) Which factor do they all share in common?

OpenStudy (mathstudent55):

Then look at the variable parts, \(x^4\), \(x^3\), and \(x^2\). \(x^4 = \color{red}{x \times x} \times x \times x\) \(x^3 = \color{red}{x \times x} \times x \) \(x^2 = \color{red}{x \times x}\) Which variable part do all three monomials have in common?

OpenStudy (prince1342):

2 is what they all have in common

OpenStudy (mathstudent55):

Yes, as far as the numbers are concerned, the only common factor is 2. What about the variables? How many x's do they have in common?

OpenStudy (prince1342):

3

OpenStudy (prince1342):

so would it be 2x^#? @mathstudent55

OpenStudy (prince1342):

2x^3 i mean

OpenStudy (mathstudent55):

??? In each line I only see 2 x's in red. Since the last monomial only has 2 x's there are only 2 x's in common. Where do you see 3 x's in all three lines above?

OpenStudy (prince1342):

I meant 3 because there were 3 lines that held at least 2 xs

OpenStudy (mathstudent55):

Don't confuse the multiplication sign with the variable x.

OpenStudy (mathstudent55):

Ok, then what is in common? \(2 \times x \times x = 2x^2\) The common factor is \(2x^2\)

OpenStudy (prince1342):

2 and x are what is in common

OpenStudy (prince1342):

so 2x^2 is what the answer would be?

OpenStudy (mathstudent55):

Yes.

OpenStudy (prince1342):

but that isnt one of the answer choices